I have tried to solve the below equation and would to verify that I am correct in my process - thank you.
x/x+3 = 2/x-3
x(x-3)=2(x+3)
x^2-3x=3x+6
x^2-5x-6=0
(x-3)(x-2)
x-3=0, x=3
x-2=0, x=2

Answer by Alan3354(69443)

I have tried to solve the below equation and would to verify that I am correct in my process - thank you.
x/x+3 = 2/x-3
x(x-3)=2(x+3)
x^2-3x=3x+6
x^2-5x-6=0
(x-3)(x-2)
x-3=0, x=3
x-2=0, x=2
------------
Sub the values into the original eqn, see if it works.
3/(3+3) =? 2/(3-3) No, 3 doesn't.
2/(2+3) =? 2/(2-1) No again.
---------
x/x+3 = 2/x-3
x(x-3)=2(x+3)
x^2-3x=2x+6
x^2-5x-6=0
----------
(x-3)(x-2) ** = x^2 - 5x + 6 (not -6)
----------
x^2-5x-6=0
(x+1)(x-6) = 0
x = 6
x = -1
-------
x-3=0, x=3
x-2=0, x=2
------------